Express in formality the concentration of 28.4 per cent in NH3 by weight having a density of 0.808 grams per ml


1
Expert's answer
2021-02-28T06:18:14-0500

Let's assume that we have 1 L of NH3 solution of 28.4%.

If its density is 0.808 grams per ml, the mass of 1 L is 808 g or 0.808 kg.

% = 100 x m(solute) / (m(solute) + m(solvent))

m(solute)NH3 = (808 x 28.4)/100 = 229.5 g

CM = n/V

n = m/M

M (NH3) = 17.03 g/mol

n (NH3) = 229.5 / 17.03 = 13.48 mol

CM (NH3) = 13.48 / 1 =13.48 M